# Distillation Dynamics: Towards Understanding Feature-Based Distillation in Vision Transformers

### v1

@@ Abstract
While feature-based knowledge distillation has proven highly effective for compressing CNNs, these techniques unexpectedly fail when applied to Vision Transformers (ViTs), often performing worse than simple logit-based distillation. We provide the first comprehensive analysis of this phenomenon through a novel analytical framework termed as “distillation dynamics”, combining frequency spectrum analysis, information entropy metrics, and activation magnitude tracking. Our investigation reveals that ViTs exhibit a distinctive U-shaped information processing pattern: initial compression followed by expansion. We identify the root cause of negative transfer in feature distillation: a fundamental representational paradigm mismatch between teacher and student models. Through frequency-domain analysis, we show that teacher models employ distributed, high-dimensional encoding strategies in later layers that smaller student models cannot replicate due to limited channel capacity. This mismatch causes late-layer feature alignment to actively harm student performance. Our findings reveal that successful knowledge transfer in ViTs requires moving beyond naive feature mimicry to methods that respect these fundamental representational constraints, providing essential theoretical guidance for designing effective ViTs compression strategies. All source code and experimental logs are provided in the supplementary material.
@@ 1 Introduction
The core challenge lies in the fundamental architectural and representational differences between CNNs and ViTs. While sophisticated feature distillation methods (hao2022manifold; yang2024vitkd; miles2024srd; fan2024scalekd; feng2025alignkd) have shown improvements, they often rely on empirical observations rather than principled theoretical guidance. What is urgently needed is a deep, mechanistic understanding of how information flows through ViT architectures and why conventional feature distillation fails—insights that can guide the development of more effective distillation methods.
@@ 2.2 Information Entropy Analysis
While frequency analysis reveals the composition of learned features, we employ information entropy to quantify their structural complexity. This analysis is explicitly framed through the lens of the Information Bottleneck (IB) principle (tishby2000ib; tishby2015dib; hong2025cib). We use Shannon entropy as a proxy for the mutual information between the input and a given layer’s representation. In this context, low entropy signifies a compressed, highly structured representation (a tight bottleneck), whereas high entropy indicates a more uniform or expansive distribution.
@@ 2.4 Distillation Methods for Validation
Both methods combine feature losses with standard KD loss: $\mathcal{L}_{\mathrm{Total}}=\mathcal{L}_{\mathrm{KD}}+\beta\mathcal{L}_{\mathrm{Feature}}$, where $\beta$ weights the feature component. More details about method, hyperparameters and experimental settings are provided in the Appendix.
@@ 5 Distillation Evolution
To deepen our understanding of the representational mismatch identified in Section 4, we examine the temporal evolution of the student’s layer-wise Shannon entropy during training. This analysis reveals how distillation strategies shape the development of the student’s information processing signature over epochs, shedding light on the dynamic interplay between the teacher’s guidance and the student’s inherent learning trajectory. By comparing the entropy profiles at checkpoints every 30 epochs (up to 300), we uncover how feature alignment alters the emergence of the U-shaped structure, explaining the observed negative transfer from late layers.
We focus on four configurations: logits-only distillation (SoftKD), frequency alignment on the first layer (SpectralKD-First), on the last layer (SpectralKD-Last), and on both the first and last layers (SpectralKD-Both). Entropy values were computed as described in Section 2.2 on DeiT-Tiny students distilled from the CaiT-S24 teacher, using ImageNet validation samples. The evolving profiles are visualized in Figure 4, with quantitative data available in the supplementary material.
@@ 5.2 Dynamics with Late-Layer Alignment
As training progresses, the last-layer entropy gradually increases, while early-layer entropy decreases. This upward drift in the last layer implies a conflict: the feature alignment loss pulls the student’s representation toward the teacher’s distributed, high-dimensional encoding, but the student’s limited capacity and the primary classification objective push for higher entropy to enable flexible feature recombination. The resulting profile retains a U-shape but with a shallower expansion phase (last-layer entropy 2.2 bits vs. 3.3 in SoftKD), potentially limiting semantic richness and contributing to the observed performance drop ($76.83\%$). This dynamic highlights negative transfer as a tug-of-war, where late-layer constraints hinder the student’s natural progression toward the teacher’s full signature.
@@ 5.3 Early-Layer Alignment and Global Effects
Combining first- and last-layer alignment (Figure 4(d)) reveals compounded effects. The profile starts with moderate entropy and evolves to a shallow increase, with minimal U-shape. Strikingly, introducing late-layer alignment causes early-layer entropy to decrease over training, inverting the trend seen in SoftKD and SpectralKD-First. This global propagation, where last-layer constraints retroactively suppress early-layer entropy, demonstrates that feature distillation has network-wide impacts, even when targeted locally. It contradicts the expected U-shape’s high initial entropy, suggesting that dual constraints over-regularize both phases, flattening the bottleneck. Despite this, the marginal performance gain ($77.08\%$) indicates a delicate balance where complementary alignments may mitigate some representational mismatches.
@@ C.1 Frequency Alignment (SpectralKD)
Baseline (No Distillation) N/A DeiT-Tiny 72.20
@@
Table 2: Distillation results using a non-fine-tuned ViT-Base (MAE) as the teacher. Accuracy is reported on ImageNet Top-1. The bolded value indicates the best performance.
@@ C.3 Experimental Setup
D. Validation with a Non-Fine-Tuned MAE Teacher
To further validate the paper’s central conclusion that negative transfer in feature-based distillation for ViTs stems from a fundamental representational mismatch in later layers, we conduct an additional experiment under a challenging “uninformative teacher” condition. The objective was to isolate the effect of feature structure from the semantic quality of the teacher’s knowledge.
We used a ViT-Base model, pre-trained with Masked Autoencoders (MAE) but not fine-tuned on ImageNet, as the teacher. This setup ensures that the teacher’s final logits are not predictive of ImageNet classes and are effectively “random” or inaccurate for the downstream classification task. The student model was the standard DeiT-Tiny. We compared two distillation strategies against a baseline:
  1. 1.
    Logits Distillation Only.Standard knowledge distillation using the softmax outputs of the non-fine-tuned teacher.
  2. 2.
    SpectralKD (Logits + Last Layer Features). Our proposed frequency alignment method applied to the last feature map layer, combined with logits distillation.
  3. 3.
    Baseline. The DeiT-Tiny model trained from scratch without any distillation.
The results of this experiment are summarized in the Table 2.
The results provide strong evidence for our hypothesis. First, using only the inaccurate logits from the non-fine-tuned MAE teacher don’t harm the student’s performance. In fact, it results in a marginal improvement ($0.15\%$) over the baseline. This demonstrates that logit-based distillation is remarkably robust and that the student model is not negatively impacted by poor-quality “soft labels”, treating them as a form of neutral regularization.
In contrast, when we introduce feature distillation from the teacher’s last layer, the student’s accuracy drop significantly by over $5$ percentage points to $67.05\%$. This severe degradation occurred even though the teacher’s features are not from a model optimized for the classification task. This outcome isolates the cause of negative transfer: it is not the semantic content of the teacher’s features that is harmful, but their fundamental structural paradigm. The late layers of the ViT-Base model, even without fine-tuning, adopt a distributed, high-dimensional encoding strategy that the smaller DeiT-Tiny is architecturally incapable of mimicking.
This experiment powerfully demonstrates that forcing a student to align with this incompatible representational structure is actively detrimental, confirming that the “representational paradigm mismatch” is the primary driver of negative transfer in ViT feature distillation.

### v2

@@ Abstract
While feature-based knowledge distillation has proven highly effective for compressing CNNs, these techniques unexpectedly fail when applied to Vision Transformers (ViTs), often performing worse than simple logit-based distillation. We provide the first comprehensive analysis of this phenomenon through a novel analytical framework termed as “distillation dynamics”, combining frequency spectrum analysis, information entropy metrics, and activation magnitude tracking. Our investigation reveals that ViTs exhibit a distinctive U-shaped information processing pattern: initial compression followed by expansion. We identify the root cause of negative transfer in feature distillation: a fundamental representational paradigm mismatch between teacher and student models. Through frequency-domain analysis, we show that teacher models employ distributed, high-dimensional encoding strategies in later layers that smaller student models cannot replicate due to limited channel capacity. This mismatch causes late-layer feature alignment to actively harm student performance. Our findings reveal that successful knowledge transfer in ViTs requires moving beyond naive feature mimicry to methods that respect these fundamental representational constraints, providing essential theoretical guidance for designing effective ViTs compression strategies. All source code and experimental logs are provided at https://github.com/thy960112/Distillation-Dynamics.
@@ 1 Introduction
The core challenge lies in the fundamental architectural and representational differences between CNNs and ViTs. While sophisticated feature distillation methods (hao2022manifold; yang2024vitkd; miles2024srd; fan2024scalekd; feng2025alignkd) have shown improvements, they often rely on empirical observations rather than principled theoretical guidance (cheng2024layerwise). What is urgently needed is a deep, mechanistic understanding of how information flows through ViT architectures and why conventional feature distillation fails—insights that can guide the development of more effective distillation methods.
@@ 2.2 Information Entropy Analysis
While frequency analysis reveals the composition of learned features, we employ information entropy to quantify their structural complexity. This analysis is explicitly framed through the lens of the Information Bottleneck (IB) principle (tishby2000ib; tishby2015dib; hong2025cib). In this context, low entropy signifies a compressed, highly structured representation (a tight bottleneck), whereas high entropy indicates a more uniform or expansive distribution.
@@ 2.4 Distillation Methods for Validation
Both methods combine feature losses with standard KD loss: $\mathcal{L}_{\mathrm{Total}}=\mathcal{L}_{\mathrm{KD}}+\beta\mathcal{L}_{\mathrm{Feature}}$, where $\beta$ weights the feature component.
@@ 5 Distillation Evolution
To deepen our understanding of the representational mismatch identified in Section 4, we examine the temporal evolution of the student’s layer-wise Shannon entropy during training. This analysis reveals how distillation strategies shape the development of the student’s information processing signature over epochs. By comparing the entropy profiles at checkpoints every 30 epochs (up to 300), we uncover how feature alignment alters the emergence of the U-shaped structure, explaining the observed negative transfer from late layers.
We focus on four configurations: logits-only distillation (SoftKD), frequency alignment on the first layer (SpectralKD-First), on the last layer (SpectralKD-Last), and on both the first and last layers (SpectralKD-Both). Entropy values were computed as described in Section 2.2 on DeiT-Tiny students distilled from the CaiT-S24 teacher.
@@ 5.2 Dynamics with Late-Layer Alignment
As training progresses, the last-layer entropy gradually increases, while early-layer entropy decreases. This upward drift in the last layer implies a conflict: the feature alignment loss pulls the student’s representation toward the teacher’s distributed, high-dimensional encoding, but the student’s limited capacity and the primary classification objective push for higher entropy to enable flexible feature recombination. This dynamic highlights negative transfer as a tug-of-war, where late-layer constraints hinder the student’s natural progression toward the teacher’s full signature.
@@ 5.3 Early-Layer Alignment and Global Effects
Combining first- and last-layer alignment (Figure 4(d)) reveals compounded effects. The profile starts with moderate entropy and evolves to a shallow increase, with minimal U-shape. Strikingly, introducing late-layer alignment causes early-layer entropy to decrease over training. This global propagation, where last-layer constraints retroactively suppress early-layer entropy, demonstrates that feature distillation has global impacts. It contradicts the expected U-shape’s high initial entropy, suggesting that dual constraints over-regularize both phases, flattening the bottleneck. Despite this, the marginal performance gain ($77.08\%$) indicates a delicate balance where complementary alignments may mitigate some representational mismatches.
@@ 6 Conclusion
7 Acknowledgments
This work was supported by the STI 2030 Major Projects (grant 2021ZD0200403) and by the Zhejiang Provincial Natural Science Foundation of China (grant LD24F030002).
@@ C.1 Frequency Alignment (SpectralKD)
Baseline (No Distillation) N/A DeiT-Tiny 74.86
@@
Logits Distillation Only ViT-Large DeiT-Tiny 76.85
SpectralKD (Logits + Last Layer) ViT-Large DeiT-Tiny 76.58
Table 2: ImageNet Top-1 accuracy of DeiT-Tiny under different distillation settings.
@@ C.3 Experimental Setup
D. Validation with More Teachers
To further validate our claim that feature-based distillation in ViTs can induce negative transfer, we conduct an additional study using two teachers for a DeiT-Tiny student, as summarized in Table 2111The baseline result in Table 2 was recomputed with our own training setup; it thus differs numerically from the value in the first arXiv version.. Following our main setup, we first consider a ViT-Base model that is pre-trained with MAE but not fine-tuned on ImageNet. In this case, the teacher’s logits are essentially uninformative for the downstream classification task. Consistently, distilling from these noisy logits already hurts the student, reducing Top-1 accuracy from $74.86\%$ (no distillation) to $72.35\%$. When we further impose last-layer feature alignment via SpectralKD, performance drops dramatically to 67.05%. This shows that forcing the student to mimic the late-layer feature maps of an uninformative teacher amplifies negative transfer rather than providing a useful signal.
We then use a supervised ViT-Large teacher. Here, logits-based distillation substantially improves the DeiT-Tiny student, achieving $76.85\%$ Top-1 accuracy. However, adding last-layer feature-map distillation again slightly degrades performance ($76.58\%$ with SpectralKD), even though the teacher itself is highly accurate. Taken together, these results reinforce our central conclusion: late-layer feature-map distillation in ViTs tends to be detrimental, regardless of whether the teacher is weak (MAE, no fine-tuning) or strong (ViT-Large). In both cases, enforcing similarity of high-dimensional feature maps pushes the student toward a representational regime it cannot realize, leading to negative transfer.